the sum of two numbers is the opposite of two. if twice the smaller number is added to three times the larger, the sum is one. find rhe numbers.
Answer by ankor@dixie-net.com(22740)

the sum of two numbers is the opposite of two.
a + b = -2 (the opposite of +2)
:
if twice the smaller number is added to three times the larger, the sum is one.
3a + 2b = 1
Multiply the 1st equation by 2, subtract from the above equation
3a + 2b = 1
2a + 2b = -4
---------------subtraction eliminates b, find a
a = 5
Find b using the 1st equation
5 + b = -2
b = -2 - 5
b = -7
:
See if that works in the 2nd equation:
3(5) + 2(-7) = 1
15 - 14 = 1
:
find the numbers. 5 & -7